OpenStudy (anonymous):

f(x)=1/(x+2) what is f'' (x) Second derivative ?

sam (.sam.):

\[f'(x)=-\frac{1}{(x+2)^2}\]

OpenStudy (anonymous):

f(x)=(x+2)^-1 f'(x)=-(x+2)^-2 f''(x)=2(x+2)^-3

sam (.sam.):

@nbouscal yes

sam (.sam.):

You just apply the power rule here

sam (.sam.):

and chin rule

OpenStudy (anonymous):

Technically apply chain rule also, but since the derivative of x+2=dx, you can basically ignore it.

OpenStudy (anonymous):

the third derivative would be -3 (x+2)^-4 ?

OpenStudy (anonymous):

Not quite. You multiply, so it would be -3*2 (x+2)^-4, or -6(x+2)^-4